OpenStudy (anonymous):

Find P(A or B), if events A and B are overlapping. P(A) = 11/18 P(B) = 5/18 p(A and B) = 2/18

OpenStudy (anonymous):

P(A or B) = P(A) + P(B) - p(A and B) =7/9

OpenStudy (anonymous):

I think it will be P(A or B) = [P(A) + P(B)] - P(A and B)